OpenStudy (anonymous):

(sin^2(x)+cos^2(x))/(1-sin^2(x))

sam (.sam.):

Use the following trig identity \[\sin^2x+\cos^2x=1\] and \[1-\sin^2x=\cos^2x\] Simplify

OpenStudy (anonymous):

So then you would get 1/cos^2(x) but does that reduce to just sec^2(x)?